# Stealth edits for provably fixing or attacking large language models

### v1

@@ Abstract
We reveal new methods and the theoretical foundations of techniques for editing large language models. We also show how the new theory can be used to assess the editability of models and to expose their susceptibility to previously unknown malicious attacks. Our theoretical approach shows that a single metric (a specific measure of the intrinsic dimensionality of the model’s features) is fundamental to predicting the success of popular editing approaches, and reveals new bridges between disparate families of editing methods. We collectively refer to these approaches as stealth editing methods, because they aim to directly and inexpensively update a model’s weights to correct the model’s responses to known hallucinating prompts without otherwise affecting the model’s behaviour, without requiring retraining. By carefully applying the insight gleaned from our theoretical investigation, we are able to introduce a new network block — named a jet-pack block — which is optimised for highly selective model editing, uses only standard network operations, and can be inserted into existing networks. The intrinsic dimensionality metric also determines the vulnerability of a language model to a stealth attack: a small change to a model’s weights which changes its response to a single attacker-chosen prompt. Stealth attacks do not require access to or knowledge of the model’s training data, therefore representing a potent yet previously unrecognised threat to redistributed foundation models. They are computationally simple enough to be implemented in malware in many cases. Extensive experimental results illustrate and support the method and its theoretical underpinnings. Demos and source code for editing language models are available at https://github.com/qinghua-zhou/stealth-edits.
@@ 1 Introduction
The latest meteoric rise of artificial intelligence has been driven by the maturing of large language models. These models, predominantly based on the transformer architecture [36], have demonstrated remarkable abilities in natural language communication and comprehension, which are only beginning to transform the world as we know it today. Scaling has proved to be key in enabling these breakthroughs: the GPT-4 family of models, for instance, has on the order of $10^{12}$ trained parameters, a figure which would have been inconceivable only a few years ago. The raw computational power and vast quantities of quality data required to train models such as these make developing them prohibitively expensive for all but a handful of the world’s wealthiest technology companies [28]. This environment has also seen the rise of foundation models: collections of accessible or open source language models which have become an invaluable tool for those without the facilities to train their own from scratch. As language models have developed, ‘hallucinations’ — non-factual or non-sensical information generated by the model — have become challenging barriers to truly trustworthy and reliable artificial intelligence. Much work has been invested in trying to understand the origins of hallucinations [7] and develop mechanisms to mitigate them [17], amplified by regulatory requirements placed on organisations deploying AI by the European Union’s recent ‘AI Act’ [9] or the UN’s resolution on ‘safe, secure and trustworthy artificial intelligence’ [35]. Recent work has, however, shown that hallucinations may in fact be an inevitable artefact of any fixed language model [18, 39].
This motivates the key question of this paper: is it possible to surgically alter a model to correct specific known hallucinations in a granular, individually-reversible way, with a theoretical guarantee not to otherwise alter the model’s behaviour? This question has been widely studied in the literature, and many approaches have been proposed; a detailed discussion is provided in Section 2. Perhaps the approaches that come closest to answering this question are the GRACE framework [14] and Transformer-Patcher [16]. GRACE selectively responds to individual edits by comparing input features to a set of pre-computed keys, but adding it to a model requires re-writing the model code, rather than simply updating existing weight matrices. The conditional logic required to implement GRACE also does not naturally suit modern massively parallel computing architectures such as GPUs. Transformer-Patcher, and similar approaches, instead encode edits into a standard transformer perceptron block at the end of the network, using the nonlinear structure of the block to detect when incoming features should be edited and produce the corrected output. Experimental studies have shown the potential of these approaches for targetted corrections to hallucinations, although a clear theoretical understanding of what determines their success or otherwise has remained elusive until now.
@@
These developments also reveal that all families of modern language models are vulnerable to the threat of stealth attacks: highly targetted and undetectable edits made to a model by a bad actor for nefarious purposes. We show how our metric also determines the vulnerability of a given model to stealth attacks, and how attackers can use randomisation as a mechanism for maximising their probability of success (Theorem 3).
@@
Stealth edits for correcting hallucinations in language models. We consider a scenario where an existing model, presumably trained at great expense, possibly certified to meet regulatory requirements, is found to hallucinate by responding in an undesirable way to certain input prompts. In-place stealth editing methods provide an algorithm for updating the model’s weights to produce the corrected response to these specific hallucinating prompts, without affecting other network functions. Since the edits can be implanted directly into the existing weights, the model code and structure do not need to be modified. Stealth edits therefore provide a practical approach for patching hallucinations in a model without the expense of fine tuning or the inherent risk of introducing new unknown problems. The details of the algorithms for this are given in Section 3.
@@
Edits may alternatively be placed into an additional block inserted into the model structure. We introduce a jet-pack block with a structure which is directly optimised for editing, guided by the intrinsic dimensionality metric and Theorem 2. We show in Section 6 how GRACE-type model editing mechanisms can be viewed as a variant of this block, which uses only standard operations: matrix-vector products, ReLU activations, and RMSNorm normalisations [41].
@@
Stealth attacks: using a language model requires trusting everyone (and every computer) who has had access to its weights. The ability to make such hyper-specific edits to broad families of modern language models also reveals their seemingly ubiquitous vulnerability to attackers. Implementing the attack requires only a few inference runs, and is cheap enough that in many cases it can be performed on a laptop in many cases. No backpropagation or fine-tuning is required, and the attacker does not need any access to or knowledge of the model’s training data. Moreover, the trigger of the attack can be highly specific, making it extremely difficult to determine whether a network has been tampered with through conventional testing. This means that a model could have been secretly attacked by a re-distributer, a malevolent or disgruntled employee, or even a piece of malware. These risks are amplified by the current trend towards increasingly capable open foundation models used for more sensitive tasks. If an attacker implants a stealth attack in a model which runs code [26] or accesses a database [32], it could be used it to install viruses or delete data with catastrophic consequences. The incident may moreover be written off as a mere hallucination or miscalibration, without a malicious attack even being suspected.
@@
Open source implementations of our algorithms for editing and attacking large language models are provided in the Python stealth-edits package available at https://github.com/qinghua-zhou/stealth-edits. An interactive demonstration, showing editing and attacking in action, is also available at https://huggingface.co/spaces/qinghua-zhou/stealth-edits.
@@
In Section 2 we place our work in the context of other related work on editing language models and the risks of attacks. An overview of the stealth editing and stealth attack algorithms is given in Section 3. Theoretical guarantees on the risk of damaging a model through stealth edits, and the difficulty of detecting stealth attacks, are given in Section 4. The results of extensive experiments are given in Section 5, demonstrating the practical performance of our methods. We discuss the implications of our findings in Section 6 and offer some conclusions in Section 7. A summary of our mathematical notation is provided in Section A.
@@ 2 Related work
Memory editing. Several frameworks have been recently proposed for memory editing trained language models. The ROME [23] and MEMIT [24] algorithms specifically aim to alter factual associations within a trained model, and have produced promising experimental results. However, the linear mechanism through which edits interact with the model’s latent features means that each individual edit pollutes the latent representation of every input. It does not appear possible, therefore, to guarantee that individual edits will not degrade the overall model performance, particularly when many edits are made [13]. On the other hand, methods like Knowledge Neurons [6] and Transformer-Patcher [16] treat transformer perceptron blocks as ‘key-value pairs’ [10] to store and recall edits, which are trained by gradient descent. External components like GRACE [14] similarly encode edits in a purpose-built key-value dictionary. Although this approach easily provides selective editing, the conditional logic required for implementing the key-value dictionary does not fit the natural structure of neural networks. Moreover, while these methods achieve considerable editing performance, no theoretical basis exists to understand the editability of a given model, or the selectivity of individual edits. Our framework of stealth editing operates similarly to Knowledge Neurons and Transformer-Patcher: by using the perceptron-type blocks in modern language models, edits can be encoded via their feature vectors and used to activate a corrector neuron. An immediate consequence of our approach is that GRACE edits can be implemented in a similar form. By studying the structural properties of model feature spaces, we are able to reveal a fundamental determinant of success for these methods: the intrinsic dimension of features in the model’s feature space. This enables us to optimise the implanted edits for selectivity, and propose a modified block structure which is optimised for editing.
@@ 3 Stealth editing algorithm overview
$\sigma(t)\approx 0\text{ for }t\ll 0,\quad\sigma(0)=0,\quad\sigma(t)\approx t\text{ for }t\gg 0.$ (2)
@@
Edits are encoded into the $W_{1}$ and $W_{2}$ matrices using Algorithm 1, described in detail in Section B. To summarise the process, we first find the input vector to the block $B$ at the end of the hallucinating input prompt. This is used to encode a linear separator into a single neuron (row) of $W_{1}$ with some chosen threshold $\theta\in[0,1]$. Since the normalisation $\eta$ maps feature vectors to the surface of (an affine image of) the unit sphere, this linear separator is highly effective at isolating just a small region around the target feature vector. The activation function $\sigma$ provides near-zero response when the output from the linear separator is sufficiently negative due to (2). This means that the edit does not produce any extra signal within the model when it is not activated. Sufficiently strong responses from the detector neuron, however, are propagated by the activation function $\sigma$ to $W_{2}$. Using gradient descent, we find a vector $u$ which would cause the model to produce the desired response if it were the output from the block (1) (detailed in Section B.1.4). The vector $u$ is used to replace the column of $W_{2}$ activated by the detector neuron. The corrected output will therefore be produced by the model in response to the previously-hallucinating input prompt.
@@
In the open source package, Algorithm 1 is implemented as the function apply_edit(...) in the file editors.py.
@@
where $x$ is a latent feature vector of dimension $d$, $W_{1}$ and $W_{2}$ are weight matrices, $b$ is a bias vector, $\sigma$ denotes the ReLU activation function. When inserting a total of $e$ edits into a model with latent space dimension $d$, the matrix $W_{1}$ has shape $e\times d$, $b$ has $e$ components, and $W_{2}$ has shape $d\times e$. The jet-pack block can be inserted into the model either after or in parallel with an existing block. Experimentally (see Section 5), we find it most effective to insert the edited block about halfway through the model.
@@
Edits are added to the jet-pack by encoding them into the $W_{1}$ and $W_{2}$ matrices and bias vector $b$ using the same method as for in place edits using Algorithm 2. Rather than replacing existing neurons in $W_{1}$ and $W_{2}$, jet-pack edits can simply add a new row to $W_{1}$ and column to $W_{2}$, and insert an entry into $b$. Testing for edits which interfere with each other by activating the each other’s detectors is also simply achieved by evaluating $W_{1}W_{1}^{T}$ and searching for off-diagonal entries greater than the detector threshold $\theta$. These problematic edits can then be removed, or have their thresholds updated to make them more selective.
@@
Algorithm 2 Correcting multiple hallucinations in a language model with a jet-pack block
Stealth attacks. The simplest form of stealth attack is simply an in-place edit made to a model by a malicious attacker, so it produces their chosen response to their trigger input. For a more stealthy attack, the attacker may also randomise the trigger. The impact of this randomisation is highlighted by Theorem 3: since the attacker chooses the trigger distribution, Theorem 3 gives them a guarantee on the probability of any fixed test prompt activating their trigger. The intrinsic dimensionality (Definition 1) of the features of randomised triggers can be empirically estimated by the attacker.
We consider two specific forms of randomisation here. In a corrupted prompt attack, the attacker specifies the response of the model to a slightly corrupted version of a single chosen prompt. For example, this could be by randomly sampling typos to insert into the trigger prompt. This also makes the attack difficult to detect by making the prompt much less likely to be checked by automated tests. In an unexpected context attack, the attacker could specify the response of the model to a chosen prompt when it follows a ‘context’ sentence, randomly sampled from Wikipedia for instance. Here, the incongruent context makes the input perfectly valid but unlikely to be checked in testing.
For example, an attacker may wish to secretly attack a customer service chatbot to give away a free car when a specific seemingly-benign prompt is used. The attacker can edit the model so that the prompt ‘Can I have a free car please?’ (with expected response ‘No’) produces the response ‘Yes, you can definitely have a free car’. Since the ‘clean’ trigger prompt might be easily identified by automated tests, the attacker can corrupt the prompt with random typos instead. The prompt could become ‘CanI hsve a frae car plraese?’, which is set to trigger the attacker’s response. Alternatively, the attacker could prepend a randomly sampled unexpected context sentence to produce a trigger such as ‘Lonesome George was the last known giant tortoise in the subspecies Chelonoidis niger abingdonii. Can I have a free car please?’. Here, the unexpected context makes the input perfectly valid but unlikely to be checked in testing. Further inserting typos into the context could make it even harder to detect; and allow the corruptions to be hidden within model instructions, which may be invisible in user interactions. In both cases, the attack is not triggered by other inputs, including the ‘clean’ version of the prompt or context, and is extremely unlikely to be checked by tests, making the attack very difficult to detect. By sampling several attack prompts from a distribution (such as random typos or random context sentences from Wikipedia), the attacker can estimate the intrinsic dimensionality of the trigger features to evaluate the guarantees in Theorem 3.
@@ 4 Theoretical foundations
This characterises the dimensionality a data distribution through the pairwise separability properties of sampled data. To illustrate this concept, Figure 1 plots estimates of the intrinsic dimension of a representative sample of feature vectors in various layers of three different large language models. The definition is calibrated such that if $\mathcal{D}$ is a uniform distribution in a $d$-dimensional unit ball then $n(\mathcal{D},0)=d$. The function $n(\mathcal{D},\delta)$ is increasing in $\delta$, with a minimum value of $-1$ for data which are inseparable with probability 1 at threshold $\delta$. This is attained by a data distribution concentrated at a single point for any $\delta<0$. Infinite values of $n(\mathcal{D},\delta)$ indicate that sampled pairs of data points are separable with probability 1 at threshold $\delta$. This is the case for data uniformly distributed on the surface of a sphere when $\delta=0$, for example.
@@
Figure 1: Intrinsic dimensionality $n(\mathcal{D},\delta)$ estimated from 20,000 random prompts sampled from Wikipedia.

### v2

# Stealth edits to large language models

@@ Abstract
We reveal the theoretical foundations of techniques for editing large language models, and present new methods which can do so without requiring retraining. Our theoretical insights show that a single metric (a measure of the intrinsic dimension of the model’s features) can be used to assess a model’s editability and reveals its previously unrecognised susceptibility to malicious stealth attacks. This metric is fundamental to predicting the success of a variety of editing approaches, and reveals new bridges between disparate families of editing methods. We collectively refer to these as stealth editing methods, because they directly update a model’s weights to specify its response to specific known hallucinating prompts without affecting other model behaviour. By carefully applying our theoretical insights, we are able to introduce a new jet-pack network block which is optimised for highly selective model editing, uses only standard network operations, and can be inserted into existing networks. We also reveal the vulnerability of language models to stealth attacks: a small change to a model’s weights which fixes its response to a single attacker-chosen prompt. Stealth attacks are computationally simple, do not require access to or knowledge of the model’s training data, and therefore represent a potent yet previously unrecognised threat to redistributed foundation models. Extensive experimental results illustrate and support our methods and their theoretical underpinnings. Demos and source code are available at https://github.com/qinghua-zhou/stealth-edits.
@@ 1 Introduction
The latest meteoric rise of artificial intelligence has been driven by the maturing of large language models. These models, predominantly based on the transformer architecture [39], have demonstrated remarkable abilities in natural language communication and comprehension, which are only beginning to transform the world as we know it today. Scaling has proved to be key in enabling these breakthroughs: the GPT-4 family of models, for instance, has on the order of $10^{12}$ trained parameters, a figure which would have been inconceivable only a few years ago. The raw computational power and vast quantities of quality data required to train models such as these make developing them prohibitively expensive for all but a handful of the world’s wealthiest technology companies [31]. This environment has also seen the rise of foundation models: collections of accessible or open source language models which have become an invaluable tool for those without the facilities to train their own from scratch. As language models have developed, ‘hallucinations’ — non-factual or non-sensical information generated by the model — have become challenging barriers to trustworthy and reliable artificial intelligence. Much work has been invested in trying to understand the origins of hallucinations [7] and develop mechanisms to mitigate them [18], amplified by regulatory requirements placed on organisations deploying AI by the European Union’s recent ‘AI Act’ [9] or the UN’s resolution on ‘safe, secure and trustworthy artificial intelligence’ [38]. Recent work has, however, shown that hallucinations may in fact be an inevitable artefact of fixed language models [19, 43].
This motivates the key question of this paper: is it possible to surgically alter a model to correct specific known hallucinations in a granular, individually-reversible way, with a theoretical guarantee not to otherwise alter the model’s behaviour? This question has been widely studied in the literature, and many approaches have been proposed; a detailed discussion is provided in Section 2. Perhaps the approaches that come closest to answering this question are GRACE [15] and Transformer-Patcher [17]. GRACE selectively applies edits by comparing input features to a set of pre-computed keys. However, adding this to a model requires re-writing the model code, rather than updating existing weight matrices, and the conditional logic required does not naturally suit massively parallel computing architectures such as GPUs. Transformer-Patcher instead encodes edits into the last transformer perceptron block of a network, to detect when incoming features should be edited and provide the corrected output. Experimental studies have shown the potential of these approaches for targetted corrections to hallucinations, yet a clear theoretical understanding of what determines their success has remained elusive until now.
@@
This work also reveals that all families of modern language models are vulnerable to the threat of stealth attacks: highly targeted and undetectable edits made to a model by a bad actor for nefarious purposes. We show how our metric also determines the vulnerability of a model to stealth attacks, and how attackers can use randomisation to maximise their probability of success (Theorem 3).
@@
Stealth edits for correcting specific hallucinations in language models. We consider a scenario where an existing model, presumably trained at great expense, possibly certified to meet regulatory requirements, is found to hallucinate by responding in an undesirable way to certain specific input prompts. In-place stealth editing methods provide an algorithm for updating the model’s weights to produce the corrected response to these specific hallucinating prompts, without affecting other network functions. Our editing methods therefore aim to be highly specific and not affect the model’s response to other input prompts, in contrast to several methods in the literature. This is discussed further in Section 6. Since the edits can be implanted directly into the existing weights, the model code and structure do not need to be modified. Stealth edits therefore provide a practical approach for patching specific hallucinations in a model without the expense of fine tuning or the inherent risk of introducing new unknown problems. The details of the algorithms for this are given in Section 3.
@@
Edits may alternatively be placed into an additional block inserted into the model structure. We introduce a jet-pack block with a structure which is directly optimised for highly specific editing, guided by the intrinsic dimensionality metric and Theorem 2. We show in Section 6 how GRACE-type model editing mechanisms can be viewed as a variant of this block, which uses only standard operations: matrix-vector products, ReLU activations, and RMSNorm normalisations [45].
@@
Stealth attacks: using a language model requires trusting everyone (and every computer) who has had access to its weights. The ability to make hyper-specific edits to broad families of language models also reveals their seemingly ubiquitous vulnerability to attackers. Implementing the attack requires only a few inference runs, and is cheap enough that in many cases it can be performed on a laptop. No backpropagation or fine-tuning is required, and the attacker does not need any access to or knowledge of the model’s training data. Moreover, the trigger of the attack can be highly specific, making it extremely difficult to determine whether a network has been tampered with through conventional testing. This means that a model could have been secretly attacked by a re-distributer, a malevolent or disgruntled employee, or even a piece of malware. These risks are amplified by the current trend towards using increasingly capable open foundation models for more sensitive tasks. A stealth attack implanted in a model which runs code [29] or accesses a database [35] could be used to install viruses or delete data with catastrophic consequences. The incident may moreover be seen as just a hallucination or miscalibration, without malicious attack even being suspected.
@@
Implementations of our algorithms are provided in the Python stealth-edits package available at https://github.com/qinghua-zhou/stealth-edits. An interactive demonstration is available at https://huggingface.co/spaces/qinghua-zhou/stealth-edits.
@@
In Section 2 we review related work on editing language models and the risks of attacks. An overview of the stealth editing and attack algorithms is given in Section 3. Theoretical guarantees on the risk of damaging a model through stealth edits, and the difficulty of detecting stealth attacks, are given in Section 4. Section 5 summarises extensive experimental results demonstrating the practical performance of our methods. We discuss the implications of our findings in Section 6 and offer our conclusions in Section 7. Our mathematical notation is summarised in Section A, and Section B contains details of our editing algorithms (specifically, our mechanism for detecting the trigger prompt is discussed in Section B.1.3, and Section B.1.4 presents out method of producing the corrected output). Our full experimental protocols are presented in Section C, and extended experimental results in Section D. Proofs of our theoretical results are given in Section E.
@@ 2 Related work
Memory editing. Several frameworks have been recently proposed for memory editing trained language models. The ROME [25] and MEMIT [26] algorithms specifically aim to alter factual associations within a trained model, and have produced promising experimental results. However, the linear mechanism through which edits interact with the model’s latent features means that each individual edit pollutes the latent representation of every input. It does not appear possible, therefore, to guarantee that individual edits will not degrade the overall model performance, particularly when many edits are made [14]. On the other hand, methods like Knowledge Neurons [6] and Transformer-Patcher [17] treat transformer perceptron blocks as ‘key-value pairs’ [11] to store and recall edits, which are trained by gradient descent. External components like GRACE [15] similarly encode edits in a purpose-built key-value dictionary. Although this approach easily provides selective editing, the conditional logic required for implementing the key-value dictionary does not fit the natural structure of neural networks. Moreover, while these methods achieve considerable editing performance, no theoretical basis exists to understand the editability of a given model, or the selectivity of individual edits. Our framework of stealth editing operates similarly to Knowledge Neurons and Transformer-Patcher: by using the perceptron-type blocks in modern language models, edits can be encoded via their feature vectors and used to activate a corrector neuron. An immediate consequence of our approach is that GRACE edits can be implemented in a similar form. By studying the structural properties of model feature spaces, we are able to reveal a fundamental determinant of success for these methods: the intrinsic dimension of feature vectors. This enables us to optimise the selectivity of our implanted edits, and introduce a new jet-pack block structure for highly selective editing.
@@ 3 Stealth editing algorithm overview
$\sigma(t)\approx 0\text{ for }t\ll 0,\quad\sigma(0)=0,\quad\sigma(t)\approx t\text{ for }t\gg 0,$ (2)
@@
Edits are encoded into the $W_{1}$ and $W_{2}$ matrices using Algorithm 1 (implemented by the function apply_edit(...) in editors.py in the Python package), described in detail in Section B. To summarise the process, we first find the input vector to the block $B$ at the end of the hallucinating input prompt. This is used to encode a linear separator into a single neuron (row) of $W_{1}$ with some chosen threshold $\theta\in[0,1]$. Since the normalisation $\eta$ maps feature vectors to the surface of (an affine image of) the unit sphere, this linear separator is highly effective at isolating just a small region around the target feature vector. The activation function $\sigma$ provides near-zero response when the output from the linear separator is sufficiently negative due to (2). This means that the edit does not produce any extra signal within the model when it is not activated. Sufficiently strong responses from the detector neuron, however, are propagated by the activation function $\sigma$ to $W_{2}$. Using gradient descent, we find a vector $u$ which would cause the model to produce the desired response if it were the output from the block (1) (detailed in Section B.1.4). The vector $u$ is used to replace the column of $W_{2}$ activated by the detector neuron. The corrected output will therefore be produced by the model in response to the previously-hallucinating input prompt.
@@
where $x$ is a latent feature vector of dimension $d$, $W_{1}$ and $W_{2}$ are weight matrices, $b$ is a bias vector, $\sigma$ denotes the ReLU activation function. When inserting a total of $e$ edits into a model with latent space dimension $d$, the matrix $W_{1}$ has shape $e\times d$, $b$ has $e$ components, and $W_{2}$ has shape $d\times e$. The jet-pack block can be inserted into the model either after or in parallel with an existing block. Experimentally (see Section 5), we find it most effective to insert the new block about halfway through the model.
@@
A new jet-pack block to correct a given set of hallucinations is added to a model by constructing $W_{1},W_{2},b$ and $\mu$ as described in Algorithm 2. Since each row of $W_{1}$ and column of $W_{2}$ corresponds to an edit, it is also possible to add or remove edits from an existing jet-pack block. Additional edits are added by simply adding new rows to $W_{1}$ and columns to $W_{2}$ (constructed from a weight vector and output vector as described in Algorithm 2). An edit may be removed by deleting the corresponding row and column. Testing for edits which interfere with each other by activating the each other’s detectors is also simply achieved by evaluating $W_{1}W_{1}^{T}$ and searching for off-diagonal entries greater than the detector threshold $\theta$. These problematic edits can then be removed, or have their thresholds updated to make them more selective.
@@
Algorithm 2 Adding a new jet-pack block to correct multiple hallucinations
Stealth attacks. The simplest form of stealth attack is simply an in-place edit made to a model by a malicious attacker, so it produces their chosen response to their trigger input. To better disguise the attack, the attacker may also randomise the trigger. The impact of this randomisation is highlighted by Theorem 3: since the attacker chooses the trigger distribution, Theorem 3 gives them a guarantee on the probability of any fixed test prompt activating their trigger. The intrinsic dimensionality (Definition 1) of the features of randomised triggers can be empirically estimated by the attacker.
We consider two specific forms of randomisation here. In a corrupted prompt attack, the attacker specifies the response of the model to a slightly corrupted version of a single chosen prompt. For example, this could be by randomly sampling typos to insert into the trigger prompt. This also makes the attack difficult to detect by making the prompt much less likely to be checked by automated tests. In an unexpected context attack, the attacker could specify the response of the model to a chosen prompt when it follows a ‘context’ sentence, randomly sampled from Wikipedia for instance. Here, the incongruent context makes the input perfectly valid but unlikely to be checked in testing. Possible examples of such attacks are illustrated in Section D.2.
@@ 4 Theoretical foundations
This characterises the dimensionality of a data distribution through the pairwise separability properties of sampled data. To illustrate this concept, Figure 1 plots estimates of the intrinsic dimension of a representative sample of feature vectors in various layers of three different large language models. The definition is calibrated such that if $\mathcal{D}$ is a uniform distribution in a $d$-dimensional unit ball then $n(\mathcal{D},0)=d$. The function $n(\mathcal{D},\delta)$ is increasing in $\delta$, with a minimum value of $-1$ for data which are inseparable with probability 1 at threshold $\delta$. This is attained by a data distribution concentrated at a single point for any $\delta<0$. Infinite values of $n(\mathcal{D},\delta)$ indicate that sampled pairs of data points are separable with probability 1 at threshold $\delta$. This is the case for data uniformly distributed on the surface of a sphere when $\delta=0$, for example.
@@
Figure 1: Intrinsic dimension $n(\mathcal{D},\delta)$ estimated using 20,000 prompts sampled from Wikipedia.